OpenStudy (anonymous):

Please help! Describe how to transform the quantity of the sixth root of x to the fifth power, to the seventh powerinto an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (anonymous):

Can you write it out? I think it looks like \[(\sqrt[6]{x}^{5})^{7} \]

OpenStudy (anonymous):

yes, thats right.

OpenStudy (anonymous):

An nth root of a number is the same as that number to the (1/n), so the 6th root of x^5 is x^(5/6).

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Then you multiply the exponent by 7.= (5*7)/6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So I'm not sure how to make the exponent any more rational without moving backwards (\[\sqrt[6]{x}^{35}\]).

OpenStudy (anonymous):

so this is as far as we can go?

OpenStudy (anonymous):

Yes, pretty much.

OpenStudy (coconutjj):